ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 170, 2009

Albiac, F. and N. J. Kalton Lipschitz structure of quasi-Banach spaces

317

Alvarez, S., D. Berend, L. Birbrair and D. Girão Resonance sequences and focal decomposition

269

Amram, M., C. Ciliberto, R. Miranda and M. Teicher Braid monodromy factorization for a non-prime K3 surface branch curve

61

Arkhipov, S. and R. Bezrukavnikov Perverse sheaves on affine flags and Langlands dual group

135

Baldwin, J. T. and A. Kolesnikov Categoricity, amalgamation, and tameness

411

Berend, D. See S. Alvarez Bezrukavnikov, R. Perverse sheaves on affine flags and nilpotent

cone of the Langlands dual group

185

Bezrukavnikov, R., M. Finkelberg and V. Ostrik On tensor categories attached to cells in affine Weyl groups, III

207

Bezrukavnikov, R. See S. Arkhipov Birbrair, L. See S. Alvarez Ciliberto, C. See M. Amram Clemens, J. D. Isomorphism of subshifts is a universal countable

Borel equivalence relation

113

Cortazar, C., M. Elgueta and J. D. Rossi Nonlocal diffusion problems that approximate the heat equation with Dirichlet boundary conditions

53

Dubickas, A. An approximation property of lacunary sequences  95

Elgueta, M. See C. Cortazar Feng D.-J. Lyapunov exponents for products of matrices and multi-

fractal analysis. Part II: general matrices

355

Finkelberg, M. See R. Bezrukavnikov Friedlander, L. and M. Solomyak On the spectrum of the Dirich-

let Laplacian in a narrow strip

337

Gan, W. T. and N. Gurevich CAP representations of G_{2} and the Spin L-function of P GSp_{6}

1

Girão D. See S. Alvarez Gurevich, N. See W. T. Gan Kalton, N. J. See F. Albiac Klartag, B. and G. Kozma On the hyperplane conjecture for ran-

dom convex sets

253

Kolesnikov, A. See J. T. Baldwin Kozma, G. See B. Klartag Levin, G. Multipliers of periodic orbits of quadratic polynomials

and the parameter plane

285

Miranda, R. See M. Amram Ostrik, V. See R. Bezrukavnikov Rossi, J. D. See C. Cortazar Schultz, C. A short proof of w_{1}^{n}(Hom(C_{2r+1}, K_{n+2})) = 0 for all n

and a graph colouring theorem by Babson and Kozlov

125

Solomyak, M. See L. Friedlander Sun, B. Positivity of matrix coefficients of representations with real

infinitesimal characters

395

Sun, Z.-W. Zero-sum problems for abelian p-groups and covers of the integers by residue classes

235

Teicher, M. See M. Amram